Daily Summary

07/01/2017 1500 PDT

Current status: WHITE - FAST-LVOS is complete!

Be sure to check out the CIRES FAST-LVOS blog ( ) for new posts!!

Dear FASTers


The last IOP ended last night with one final high ozone day in Clark County to conclude the measurement phase of FAST-LVOS. The team has dispersed, with Guillaume, Scott, and Jeff packing up their equipment this morning and heading to Boulder with TOPAZ, microDop, and the mobile lab. They’re somewhere in Arizona or maybe even Utah by now. Chance flew back to Boulder yesterday, and Zaheer and Dani are off to California. I’m sitting here at McCarran waiting for my flight home. Thank you one and all for your hard work. We have amassed an amazing data set that will keep us busy for a long time to come. Thanks again to Zheng and Rodney and the CCDAQ for their support.